OpenStudy (anonymous):

What is the sum of a 22-term arithmetic sequence where the first term is 54 and the last term is -30? Select one: A. 264 B. 288 C. 312 D. 336

OpenStudy (anonymous):

the sum eq for arith is Sn=n(t1+t2)/2 n= how many terms there are (22) t1= first term tn= last term it's basically a plug and chug problem

OpenStudy (anonymous):

I don't know how that works

OpenStudy (anonymous):

So you put in Sn= 22(54+ -30)/2 Simplify... Sn= 22(24)/2 The answer... Sn= 264